Accelerator Update for May 11, 2009 - May 13, 2009 |
Monday May 11 The day shift began with Operations monitoring the Tevatron (TeV) store 7039 (164.31E30) and the Recycler stash (143.77E10), with the Antiproton Source stacking (21.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. Operations disabled beam to Pbar, MTest, NuMI, and MiniBooNE at 12:10 PM. This was to allow Linac personnel the opportunity to replace a driver and switch tube for a Linac RF station (LRF4). Linac personnel brought LRF4 back online at 4 PM and Linac resumed sending beam. During the LRF4 repairs, the LRF5 water system nitrogen fill line repairs were completed, Booster personnel accessed their tunnel for RF repairs, and M01 and M02 were accessed for sump water samples. A Recycler kicker (R:KPS2A) tripped off with a cooling fluid (fluorinert) low-level indication; the system was checked and jumpered. Operations disabled the MTest beam at 6 PM. Operations terminated store 7039 at 9:53 PM and put the TeV into HEP shot setup. Operations established store 7040 at 11:30 PM with an average initial luminosity of 323.04E30. Operators prepared MI for access. Operations transferred antiprotons to the Recycler as appropriate. Tuesday May 12 The midnight shift began with Operations monitoring store 7040 (269.08E30) and the stack (43.1E10), and with the Recycler, NuMI, and MiniBooNE off due to the MI access. Kicker personnel had R:KPS2A repaired by 2 AM. Operations resumed sending beam to MiniBooNE at 2:24 AM, followed shortly by beam to Pbar and NuMI. The Pbar beam did not last long. A septum (D:ISEPV) started tripping off, first on cap bank overvoltage and then on a ground fault. Operations called in Pbar personnel to investigate the problem. Operations loaded a NuMI-only timeline into the Time Line Generator (TLG). Operators prepared the Pbar rings for access to investigate a possible bad magnet for D:ISEPV. There was some downtime caused by LRF4 tripping off with a power amplifier cathode overvoltage fault. MIRF17 tripped off on a choke over temperature indication. The station would not run for more than a couple of beam pulses before tripping off again, so it was left off and bypassed. Operations established beam to MTest around 4 AM. The Pbar Debuncher D:ISEPV ground fault investigation lasted most of the day shift; the septum recovered enough to allow stacking. There was downtime due to various Linac RF station problems. Operations disabled the MTest beam at 6 PM. Operations terminated store 7040 at 6:03 PM. TeV power supply work and accesses into both CDF and D0 followed. Operations put the TeV through dry squeezes and wet squeezes and then put the TeV into shot setup. Operations transferred antiprotons to the Recycler as appropriate. Wednesday May 13 The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (249.6E10), with Pbar stacking (23.7E10), and with NuMI and MiniBooNE taking beam. Operations established store 7044 at 1:05 AM with an average initial luminosity of 275.31E30. Operations established beam to MTest at 4 AM. Operations transferred antiprotons to the Recycler as appropriate. The Plans for Wednesday The plans for today are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 05/13/2009 email Fermilab |
| Security, Privacy, Legal | |